OpenStudy (anonymous):

Assume that y varies inversely with x. If y = 8.8 when x = 2.4, find x when y = 2.1. Round to the nearest tenth

OpenStudy (jim766):

here is the equation y varies inversely to x \[y = \frac{ K }{ x }\]

OpenStudy (jim766):

now substitute in what they gave you \[8.8=\frac{ k }{ 2.4 }\] can you tell me what k =

OpenStudy (jim766):

multiply 2.4 on both sides

OpenStudy (anonymous):

isnt it 1.6

OpenStudy (jim766):

8.8 * 2.4 = 21.12

OpenStudy (jim766):

k = 21.12 do you see where I get that?

OpenStudy (anonymous):

yes

OpenStudy (jim766):

now put the k you just figure out, in with the 2nd part the gave you.. \[y =\frac{ k }{ x }\] \[2.1=\frac{ 21.12 }{ x }\] , now solve for x

OpenStudy (jim766):

when you mult x on both sides you would get 2.1x = 21.2 now divide both sides by 2.1

OpenStudy (anonymous):

its 10.1

OpenStudy (jim766):

yes, does it make sense?